# PinnDE: Physics-Informed Neural Networks for Solving Differential Equations

### v1

@@ 3 PinnDE
Physics-informed neural networks for differential equations (PinnDE) is an open-source library in Python providing the ability to solve differential equations with both physics-informed neural networks (PINNs) and deep operator networks (DeepONets). PinnDE is built in mainly using TensorFlow tensorflow2015-whitepaper with a small amount of JAX jax2018github , serving as the backend, two popular software packages for deep learning. Functionality to provide multiple backends and other deep learning frameworks such as PyTorch for use with PinnDE is in development and will be added in a future version.
@@
The goal of PinnDE is to provide a user-friendly interface for solving systems of differential equations which can easily be shared with collaborators even if they do not have experience working with machine learning in Python. Many alternative packages for solving differential equations with PINNs or DeepONets, such as DeepXDE lu2021deepxde and PINA coscia2023physics , are powerful packages in this field. They provide a high level of customization and control in how a specific problem is being solved. This naturally leads to a higher amount of software needing to be written and understood by a user, which when sharing with collaborators can lead to difficulties if not everyone understands the package to the same degree. PinnDE is a tool in which we aim to bridge this gap between functionality and ease-of-use, stripping away complexity from the code resulting in an easy to read interface non-specialists can understand. This package has user set conditions defining the type of differential equations, initial conditions, boundary conditions, and model parameters, which is then fully built with a specific training routine already implemented based off of theses flags. This allows users to have a small amount of interfacing while still being able to have access to a wide range of applications without needing to do any low-level manual implementations.
@@ 3.1 Overview
Figure 1 shows the flowchart of how a user goes about solving a system of differential equations using PinnDE. Note how a user only interfaces directly with Boundaries, Initials and Solvers. This gives a user a small amount of requirements to specify, all of which are related to the problem specification itself as well as to the basic neural network architecture and optimization procedure, but as will be demonstrated below still does not take away from the users’ ability to interact with all further parts of the solution process.
Figure 1: Flowchart of the package structure of PinnDE. Black lines represent forward function calls, red lines represent returned values.
A user first initializes the boundaries of their problem using a single function from the Boundaries module. If the problem is time-dependent and involves initial conditions, a user initializes the initial conditions using a single function from the Initials module. Then a user calls a single function in one of the Solvers modules corresponding to the system of differential equations they wish to solve. This function call creates a solution object corresponding to the Solution Object modules. This class in its constructor sets up the corresponding Collocation Points and Models to generate collocation points and a neural network with specifications from user in Solvers call. A user then calls train_model() on the returned object class invokes a corresponding Training Selection function which directs the problem and generated data to a Specific Training File, where the model is trained and all training data is returned to the user through the Solution Object from the original Solvers call.
In these next sections we give a brief overview of what boundaries, solvers, and which models are currently available in PinnDE, specifying limitations in implementation if possible. We only outline the general ability of PinnDE, a full API and tutorials can be found at https://pinnde.readthedocs.io/en/latest/.
@@ 3.1.1 Initial and Boundaries
The boundaries module provides boundary generation functions for partial differential equations. Ordinary differential equations can be solved with both initial and boundary values, however the values are passed directly into the solvers function, resulting in no need of a function call from any separate module. PinnDE is presently designed to solve equations on rectangular domains. The boundaries currently available to be used are Periodic, Dirichlet, and Neumann for both equations of variables $(t,x_{1})=(t,x)$ and of $(x_{1},x_{2})=(x,y)$. That is, PinnDE currently supports (1+1)-dimensional evolution equations or two-dimensional boundary value problems. Higher-dimensional problems will be added in a future version.
PinnDE also offers both soft and hard constraining for these boundaries. The availability of what PinnDE provides is further described in Section 3.1.3 where the models of PinnDE are described. Here we give some examples of specific implementations of how hard constraining of a boundary (temporal or spatial) is done in PinnDE.
For constraining a periodic boundary condition on a PDE in variables $(t,x)$ over a spatio-temporal domain $[t_{\rm 0},t_{\rm f}]\times\Omega$, $\Omega=[x_{\rm l},x_{\rm r}],x_{\rm l},x_{\rm r}\in\mathbb{R}$, we only have to hard-constrain the initial conditions as periodic boundaries can be implemented using a coordinate transform layer in the neural network itself, by using the transformation $f(x)=(\cos 2\pi x/(x_{\rm r}-x_{\rm l}),\sin 2\pi x/(x_{\rm r}-x_{\rm l}))$ applied to the spatial input $x$ to the neural network, see bihlo2022physics for further details. This ensures that the solution will be periodic on the interval $[x_{\rm l},x_{\rm r}]$.
As an example for hard constraining the initial condition itself, we assume that the equation is of second order in $t$. Then, we can include the initial condition as a hard constraint in the neural network surrogate solution $u^{\theta}(t,x)$ using the ansatz
@@
Where $u_{0}$ and $\frac{\partial u}{\partial t}\big{|}_{t=t_{\rm 0}}$ represent the initial conditions at $t=t_{\rm 0}$, and $\mathcal{N}^{\theta}(t,x)$ is the neural network output. Multiple variations of possible hard constraining formulas for this problem can be created which enforces the desired initial conditions. We found this particular equation to be effective for many problems we have considered, and it is therefore what is implemented in PinnDE as a default method. Alternative hard constraints will be implemented in a future version of PinnDE.
We also present the hard constraint for a Dirichlet boundary-value problem for a PDE in variables $(x,y)$ over the rectangular domain $\Omega=[x_{\rm l},x_{\rm r}]\times[y_{\rm l},y_{\rm u}]$, with $x_{\rm l},x_{\rm r},y_{\rm l},y_{\rm u}\in\mathbb{R}$. The approximate solution $u^{\theta}(x,y)$ is given by the equation
@@
This formula enforces the specific boundary conditions for all values of the neural network. This is based on what is described in lagaris1998artificial and is what is implemented in PinnDE.
3.1.2 Solvers
The solvers are split into ode_Solvers and pde_Solvers, where each provide different functions for solving specific types of problems. The functions currently available are:
Table 1: Functions for solving PDEs in PinnDE
PINNs DeepONets
solvePDE_tx() solvePDE_DeepONet_tx()
solvePDE_xy() solvePDE_DeepONet_xy()
Table 2: Functions for solving ODEs in PinnDE
PINNs DeepONets
solveODE_IVP() solveODE_DeepONet_IVP()
solveODE_BVP() solveODE_DeepONet_BVP()
solveODE_System_IVP() solveODE_DeepONetSystem_IVP()
@@ 3.1.3 Models
PinnDE currently provides the ability to use PINNs and DeepONets for solving systems of differential equations. The number of layers and nodes can both be selected within each solver function, as well as either soft or hard constraining of the network. We employ multi-layer perceptron for all neural network solution surrogates, and implement DeepONets using the architecture proposed in lu2019deeponet . Both physics-informed neural networks and physics-informed deep operator networks are trained by minimizing the composite loss function (2) to approximate the solution (or solution operator) of a system of differential equations. Derivative approximations are computed using automatic differentiation baydin2018automatic . Longer time integrations using DeepONets via time-stepping is realized using the procedure of wang2023long , which iteratively applies the learned solution operator to its own output to advance the solution for arbitrary time intervals, similar as done in standard numerical methods.
Hard constraining is done using and expanded upon from the ideas from lagaris1998artificial and brecht2023improving , which outline basic equations to force a network output to conform to initial or boundary constraints, explained further in Sections 2.1 and 3.1.1. However currently not all boundary–equation type combinations have a hard constrained model designed for, but this is can be implemented via user’s requests if the need for them arises. We outline in Tables 3 and 4 what does and what does not currently have the ability to be used in terms of hard constraints.
Table 3: Hard constraints for initial and boundary value problems available for solving ordinary differential equations in PinnDE.
ODE Problems/Models PINN DeepONet
IVP Soft, Hard Soft, Hard
BVP Soft, Hard Soft, Hard
System of IVPs Soft Soft
Table 4: Hard constraints for initial and boundary value problems available for solving partial differential equations in PinnDE.
PDE Problems/Models PINN DeepONet
$(t,x)$ Periodic Boundaries Soft, Hard Soft, Hard
$(t,x)$ Dirichlet Boundaries Soft Soft
$(t,x)$ Neumann Boundaries Soft Soft
$(x,y)$ Periodic Boundaries Soft, Hard Soft, Hard
$(x,y)$ Dirichlet Boundaries Soft, Hard Soft, Hard
$(x,y)$ Neumann Boundaries Soft Soft
@@ 3.2 Usage
Step 1 Define the domain for the independent variables
Step 2 Define the number of collocation points along each boundary to use
Step 3 Define boundary functions as Python lambda functions,
and use pde_Boundaries_2var to set up Boundaries
@@
initial value points, and use pde_Initials to set up Initials
Step 5 Define the equation as a string, number of collocation points for the pdes,
and number of training epochs
@@
Step 7 Call corresponding solvePDE function from Solvers and input data to solve equation
Step 8 Call train_model(epochs) on returned object from solvePDE to train the network
Step 9 Use returned class to plot data with given functions or take data from class
@@ 4 Examples
In this section we present a few short examples of how PinnDE can be used to solve a system of ordinary differential equations, and multiple partial differential equations using different methods implemented in PinnDE. We provide a concise description of each equation and a comparison of each trained model against the analytical solution using the mean squared error as comparison metric. All code for these examples is provided in Appendix A.
4.1 System of ordinary differential equations
@@
We first solve a system of ordinary differential equations with a deep operator network which uses soft constrained initial conditions. We solve the system of second order ordinary differential equations
$u^{\prime\prime}(t)+u(t)=0,\quad v^{\prime}(t)+u(t)=0,$ (7a)
with initial conditions
$u(0)=0.5,\quad u^{\prime}(0)=1,\quad v(0)=2.$ (7b)
We solve this initial value problem over the interval $t\in[0,1]$ using a deep operator network.
@@
Figure 2 presents the numerical solution and the exact solution as well as the time series of the loss for training the network. The analytical solution for this problem is $u(t)=-0.5\sin t+\cos t+1$ and $v(t)=\sin t+0.5\cos t$.
@@
Figure 2: Solution predicted using a DeepONet against the exact solution for system (7) (left) and the time series of the physics-informed loss over the training period of the DeepONet (right).
@@
The specific architecture used was a DeepONet with 4 hidden layers with 40 nodes per layer, using the hyperbolic tangent as activation function, and Adam as the chosen optimizer. We use 150 collocation points across the domain for the network to learn the solution, sampled with Latin hypercube sampling. We sample 5000 different initial conditions for the DeepONet in the range of $v(0),u(0),u^{\prime}(0)\in[-3,3]$. We train this network for 20000 epochs which is where the epoch loss starts to level out. For the following examples, we will also train until a minimum loss has been reached.
@@
We further demonstrate the time-stepping ability using a DeepONet in PinnDE. We time-step the trained DeepONet for 10 steps, increasing the domain to $[0,10]$, and give the error of the neural network solution against the exact solution. The learned neural network solution operator demonstrates the ability to time-step with a consistently low error.
@@
Figure 3: Time-stepped solution predicted by the trained DeepONet against the exact solution for system (7) (left), and the error between the network and analytical solution over the time-stepped solution interval (right).

### v2

@@ 3 Physics-informed neural networks for differential equations (PinnDE)
Physics-informed neural networks for differential equations (PinnDE) is an open-source library in Python providing the ability to solve differential equations with both physics-informed neural networks (PINNs) and deep operator networks (DeepONets). PinnDE is built using TensorFlow tensorflow2015-whitepaper and JAX jax2018github , serving as the backend, two popular software packages for deep learning. Functionality to provide multiple backend and other deep learning frameworks such as PyTorch for use with PinnDE is in development and will be added in a future version.
@@
The goal of PinnDE is to provide a user-friendly interface for solving systems of differential equations which can be easily shared with collaborators even if they do not have experience working with machine learning in Python. Many alternative packages for solving differential equations with PINNs or DeepONets, such as DeepXDE lu2021deepxde and PINA coscia2023physics , are powerful packages in this field. They provide a high level of customization and control in how a specific problem is being solved. This naturally leads to a higher amount of software needing to be written and understood by a user, which when sharing with collaborators can lead to difficulties if not everyone understands the package to the same degree. PinnDE is a tool in which we aim to bridge this gap between functionality and ease-of-use, stripping away complexity from the code resulting in an easy to read interface non-specialists can understand. This package has users create small, modular objects which represent components of a differential equation problem, which then requires just minimal interfacing with the package albeit still being able to have access to a wide range of applications, without the need to do any low-level manual implementations.
@@ 3.1 Overview
Figure 1 shows the data flow of main modules involved in solving a system of differential equations using PinnDE. Each individual module gives the user a small amount of requirements to specify, all of which are related to the problem specification itself as well as to the basic neural network architecture and optimization procedure, but as will be demonstrated below still does not take away from the users’ ability to interact with all further parts of the solution process.
Figure 1: Flow chart of the module structure of PinnDE
A user first initializes the domain of their problem using a single function from the Domain module. Next, the boundary conditions of the problem are initialized with the Boundaries module. If the problem is time-dependent and involves initial conditions, the initial conditions are defined using a single function from the Initials module. Then a single function in the Data module is being called corresponding to the system of differential equations one wishes to solve. This function call creates a data object, which is what is passed to a function call in the Model module to create the desired model to solve the specified differential equations. As a next step one calls the train() on one’s model, invoking the training routine for the specific problem with desired functionality (hard constraining, adaptive collocation points, meta-learned adaptivity, etc). Lastly, PinnDE then provides a plotting module to quickly plot losses and visualization of results on specific dimension equations.
In these next sections we give a brief overview of the domains, boundaries, data, which models and what adaptive point sampling strategies are currently available in PinnDE, specifying limitations in implementation if applicable. We only outline the general ability of PinnDE, a full API and tutorials can be found at https://pinnde.readthedocs.io/en/latest/.
3.1.1 Domains
The domain module is the first which must be initialized when using PinnDE. Currently we provide two general types of domains implemented, $d$ dimensional hyper-rectangles and $d$ dimensional ellipsoids. Both of these domains allow specifying a domain in $d$ dimensions, denoted in PinnDE as $x1,\dots,xd$. Both of these have a spatial domain function, as well as a temporal domain counterpart, which is used when defining a time dependent system of partial differential equations, which allows the time dimension, denoted $t$ in PinnDE, to be used. This enables solving $(1+d)$-dimensional evolution equations or $d$-dimensional boundary value problems. We also provide a general API for boundary functions with the abstract boundaries class, which when implemented allow users to define custom-domains which will work with the PinnDE architecture.
@@ 3.1.2 Initial and Boundaries
The Boundaries module provides boundary generation functions for ordinary and partial differential equations. The boundaries currently available to be used for partial differential equations are Periodic, Dirichlet, and Neumann. We further provide the boundary ODEicbc, which is designed to be used to specify initial or boundary conditions for systems of ordinary differential equations. All boundary functions require a domain to be passed in for which the boundary is being defined. These boundaries can be specified for any $(1+d)$-dimensional system of evolution equations or $d$-dimensional boundary value problems. Specific implementation of the boundary values depends on the domain being described.
PinnDE also offers both soft and hard constraining for a selection of these boundaries in specific dimensions. The availability of what PinnDE provides is further described in Section 3.1.4 where the Models of PinnDE are described. Here we give some examples of specific implementations of how hard constraining of a boundary (temporal or spatial) is done in PinnDE.
Periodic boundary conditions over $n$ spatial dimensions $(x_{1},\dots,x_{n})$ of a differential equation over the hyper-rectangle $(x_{1},\dots,x_{n})\in[x_{\rm l_{1}},x_{\rm r_{1}}]\times\dots\times[x_{\rm l_{n}},x_{\rm r_{n}}]$ are enforced using a coordinate transform layer in the neural network itself, by using the transform $f(x_{i})=(\cos 2\pi x_{i}/(x_{\rm r_{i}}-x_{\rm l_{i}}),\sin 2\pi x_{i}/(x_{\rm r_{i}}-x_{\rm l_{i}}))$ applied to each spatial input $x_{i}$. See bihlo2022physics for further details.
For constraining a periodic boundary condition of a $(1+1)$-dimensional PDE in variables $(t,x)$ over a spatio-temporal domain $t_{\rm 0},t_{\rm f}]\times\Omega$, $\Omega=[x_{\rm l},x_{\rm r}],x_{\rm l},x_{\rm r}\in\mathbb{R}$, we only have to hard-constrain the initial conditions as periodic boundaries can be enforced exactly using a coordinate transform layer described above. For hard constraining the initial condition itself, we assume that the equation is of second order in $t$. Then, we can include the initial condition as a hard constraint in the neural network surrogate solution $u^{\theta}(t,x)$ using the ansatz
@@
Where $u_{0}$ and $\frac{\partial u}{\partial t}\big{|}_{t=t_{\rm 0}}$ represent the initial conditions at $t=t_{\rm 0}$, and $\mathcal{N}^{\theta}(t,x)$ is the neural network output. Multiple variations of possible hard constraining formulas for this problem can be created which enforce the desired initial conditions. We found this particular hard constraint ansatz to be effective for many problems we have considered, and it is therefore used in PinnDE as a default hard constraint. However more hard constraints are available in PinnDE, including custom hard constraints.
We also present the hard constraint for a two-dimensional Dirichlet boundary-value problem for a PDE in variables $(x,y)$ over the rectangular domain $\Omega=[x_{\rm l},x_{\rm r}]\times[y_{\rm l},y_{\rm u}]$, with $x_{\rm l},x_{\rm r},y_{\rm l},y_{\rm u}\in\mathbb{R}$. The approximate solution $u^{\theta}(x,y)$ is given by the equation
@@
This formula enforces the specific boundary conditions for all output values of the neural network. This is based on what is described in lagaris1998artificial and is the default hard constraint for two-dimensional spatial problems over a rectangular domain in PinnDE. Other alternative hard constraints do exist in PinnDE if applicable to the scenario, and the ability for user’s to create custom hard constraints also exist.
Hard constraints can be designed for general domains, however the particular domain features impact the hard constraining function, making them cumbersome to design. In PinnDE, we give designed hard constraints pre-made for the most popular use cases of domain-boundary combinations. We currently provide built-in hard constraints for rectangular domains. We provide hard constraints in the following cases: Single ODEs, single $(1+1)$-dimensional PDEs with periodic boundaries, and single $2$-dimensional PDEs with periodic and Dirichlet boundaries. However, PinnDE is built for custom hard constraints to be able to be created and added to the network, allowing for any user-specific hard constraint to be implemented on the network. In doing this, we cover functionality for most use cases while still having the possibility for customization for any problem.
3.1.3 Data
The Data module of PinnDE combines and creates sampled collocation points for training. Each model has its own specific data object, with a time data counterpart to incorporate a time dimension when training for spatio-temporal equations. The data module always takes the Domain and Boundaries as arguments. If using a time domain, it also must take in an Initials. Then model specific information may be passed depending on the specific data being created. For example, DeepONet data objects need the number of sensor points to sample, and inverse PINN data objects need the sampled data over the problem domain and all functions being solved for. This is then the final step before defining the model.
@@ 3.1.4 Models
PinnDE currently provides the ability to use PINNs and DeepONets for solving systems of differential equations. For PINNs, we provide standard forward PINNs as well as inverse PINNs for solving inverse problems. The number of layers and nodes can both be selected within each model function. The outer and inner activation function used within each model can be chosen by the user as well. The user also has the choice to either soft or hard constrain the network. We employ multi-layer perceptron for all neural network solution surrogates, and implement DeepONets using the architecture proposed in lu2019deeponet . Both physics-informed neural networks and physics-informed deep operator networks are trained by minimizing the composite loss function (2) to approximate the solution (or solution operator) of a system of differential equations. Derivative approximations are computed using automatic differentiation baydin2018automatic . Longer time integrations using DeepONets via time-stepping is realized using the procedure of wang2023long , which iteratively applies the learned solution operator to its own output to advance the solution for arbitrary time intervals, similar as done in standard numerical methods.
3.1.5 Adaptive
When calling train() on a model from PinnDE, the user is allowed to pass in an adaptive point sampling strategy created from the Adaptive’s module. We currently provide three different adaptive sampling strategies which can be employed without needing to do any manual implementation: residual-based adaptive refinement, residual-based adaptive distribution, and residual-based adaptive refinement with distribution (RAR, RAD, RAR-D). These strategies are based on the research reported in lu2021deepxde ; wu2023comprehensive , and were shortly summarized in Section 2.4. These strategies can be used with both forward and inverse PINNs, and can also be applied in combination with L-BFGS optimization.
@@ 3.2 Usage
Step 1 Use the Domain module to define the domain for the independent variables
Step 2 Define the boundary functions as Python lambda functions,
as well as number of sample points to use on boundary
Step 3 Use the Boundaries module to set up the desired boundary
@@
initial value points, and use the Initials module to set up initial conditions
Step 5 Define the Data object corresponding to the model being used
@@
Step 7 Define the equation to solve, and use Model module and input data to solve equation
Step 8 Call train(epochs) on returned model to train the network
Step 9 Use trained model to plot data with given functions or take data from class
@@ 4 Examples
In this section we present a few short examples of how PinnDE can be used to solve systems of ordinary differential equations, and systems of partial differential equations using the different methods implemented in PinnDE. We provide a concise description of each problem and a comparison of each trained model against the analytical solution or a numerical reference solution obtained using a standard numerical integrator using the mean squared error as comparison metric. All code for these examples is provided in Appendix A.
4.1 Rigid body
@@
We first consider the Euler equations for the free rigid body holm2009geometric , a system of ordinary differential equations, with a physics-informed network which uses soft constrained initial conditions. Specifically, we solve the system of order ordinary differential equations in angular momentum representation
$\frac{dw_{1}}{dt}$ $=\frac{I_{2}-I_{3}}{I_{2}I_{3}}w_{2}w_{3},$ (10a)
$\frac{dw_{2}}{dt}$ $=\frac{I_{3}-I_{1}}{I_{1}I_{3}}w_{1}w_{3},$ (10b)
$\frac{dw_{3}}{dt}$ $=\frac{I_{1}-I_{2}}{I_{1}I_{2}}w_{1}w_{2},$ (10c)
where $w=(w_{1},w_{2},w_{3})$ is the angular momentum of the rigid body, and $I_{1},I_{2},I_{3}$ are the principle moments of inertia. For this benchmark we use $I_{1}=0.2,I_{2}=0.3,I_{3}=0.4$, and the initial condition $w=(1,1,1)$. We solve this initial value problem over the interval $t\in[0,2.5]$.
@@
Figure 2 presents the neural network solution against a numerical solution as well as the time series of the loss for training the network. The standard numerical reference solution for this problem was computed using the standard fourth-order Runge–Kutta method.
@@
Figure 2: Solution predicted using a PINN against the numerical solution for the rigid body system (10) (left), the difference of neural network solution and numerical solution (middle), and the time series of loss from training the network (right).
@@
The specific architecture used was a PINN with 4 hidden layers with 60 nodes per layer, using the hyperbolic tangent as activation function, and Adam and L-BFGS for fine-tuning was chosen as the optimizers. We use 3000 collocation points across the domain for the network to learn the solution, sampled with Latin hypercube sampling. We train this network for 20000 epochs using Adam, then 5000 iterations using L-BFGS, implementing a two-step training routine. The start of L-BFGS optimization during training is highlighted in Fig. 2. We stop training when the epoch loss converges towards the solution. For the following examples, we will also train until a minimum loss has been reached.
The code for solving this system of equations can be found in Appendix A.1.
4.2 Lorenz model
We next use PinnDE to solve the Lorenz–1960 model, a system of ordinary differential equations, with a deep operator network which uses soft-constrained initial conditions. The dynamical system is
$\frac{dx}{dt}$ $=kl(\frac{1}{k^{2}+l^{2}}-\frac{1}{k^{2}})yz,$ (11a)
$\frac{dy}{dt}$ $=kl(\frac{1}{l^{2}}-\frac{1}{k^{2}+l^{2}})xz,$ (11b)
$\frac{dz}{dt}$ $=\frac{kl}{2}(\frac{1}{k^{2}}-\frac{1}{l^{2}})xy,$ (11c)
where in this benchmark we chose $k=2,l=1$, and we choose initial conditions $x(0)=0.5,y(0)=0.75,z(0)=1$. We solve this initial value problem over the interval $t\in[0,1]$.
Figure 3 presents the neural network solution against a numerical solution as well as the time series of the loss for training the network. The standard numerical reference solution for this problem was computed using the Runge–Kutta method.
Figure 3: Solution predicted using a DeepONet against the numerical solution for the Lorenz system (11) (left), the difference of neural network solution and numerical solution (middle), and the time series of loss from training the network (right).
The specific architecture used was a DeepONet with both branch net and trunk net being multi-layer perceptrons having 4 hidden layers with 40 nodes per layer, using the hyperbolic tangent as activation function, and Adam as the chosen optimizer. We use 3000 collocation points across the domain for the network to learn the solution, sampled with Latin hypercube sampling. We sample 5000 different initial conditions for the DeepONet’s branch network in the range of $x(0),y(0),z(0)\in[-2,2]$. We train this network for 75000 epochs using Adam. This is when the epoch loss converges towards its minimum.
@@ 4.2 Lorenz model
We further demonstrate the time-stepping ability using a DeepONet in PinnDE in Fig. 4. We time-step the trained DeepONet for 6 steps, increasing the temporal domain to $[0,6]$, and give the error of the neural network solution against the corresponding Runge–Kutta numerical solution. The learned neural network solution operator demonstrates the ability to time-step with a consistently low error.
@@
Figure 4: Time-stepped solution predicted by the trained DeepONet against the numerical reference solution obtained using a fourth-order Runge–Kutta method for system (11) (left), the difference of neural network solution and numerical solution (right).
The code for solving this system of equations can be found in Appendix A.2.
4.3 Dam break problem
We next use PinnDE to solve the classical one-dimensional dam break problem with flat bottom topography for the shallow-water equations. We use soft-constrained initial conditions and Dirichlet boundary conditions. Specifically, we solve the one-dimensional shallow water equations given by
$\frac{\partial h}{\partial t}+\frac{\partial(hu)}{\partial x}=0,$ (12a)
$\frac{\partial(hu)}{\partial t}+\frac{\partial}{\partial x}\left(hu^{2}+\frac{1}{2}gh^{2}\right)=0,$ (12b)
where $h(t,x)$ represents the fluid depth and $u(t,x)$ represents the fluid velocity. The analytical solution for this problem is given by
$gh_{a}(t,x)=\begin{cases}a_{0}^{2}&\text{if }x-x_{0}<-a_{0}t\\ \frac{1}{9}(2a_{0}-(x-x_{0})/t)^{2}&\text{if }-a_{0}t<x-x_{0}<2a_{0}t\\ 0&\text{if }x-x_{0}>2a_{0}t\end{cases}$
$u_{a}(t,x)=\begin{cases}0&\text{if }x-x_{0}<-a_{0}t\\ \frac{2}{3}(a_{0}+(x-x_{0})/t)^{2}&\text{if }-a_{0}t<x-x_{0}<2a_{0}t\\ 0&\text{if }x-x_{0}>2a_{0}t\end{cases}$
where $a_{0}=\sqrt{gh_{0}}$, see bokhove2005flooding for further details. This analytical solution gives the initial and boundary conditions. In this benchmark, we use initial parameters $h_{0}=0.75,x_{0}=-0.25$, and we set $g=1$. We consider this benchmark over the spatio-temporal domain $(t,x)\in[0.01,1]\times[-\pi,\pi]$. We compare the neural network solution to the analytical solution for both $h(t,x)$ and $h(t,x)u(t,x)$ in Figure 5.
Figure 5: Solution predictions using a PINN for the dam break problem (12) (left), the analytical solution (middle), and the difference between them (right). Top: Water height $h(t,x)$. Bottom: Momentum $h(t,x)u(t,x)$.
For this example we used a PINN with 4 hidden layers with 60 nodes per layer, the hyperbolic tangent activation function, and trained the neural network using the Adam optimizer. A total of 1000 initial value collocation points were used for the network to learn the initial condition, 2000 boundary collocation points, and 15000 collocation points were used across the spatio-temporal domain for the network to learn the solution, both sampled using Latin hypercube sampling. The minimum of the loss was reached after 35000 epochs of training.